OpenStudy (anonymous):

solve for x 4x(x-8)-5x(x-7)=-4

OpenStudy (anonymous):

x = -1, 4

OpenStudy (anonymous):

it comes out to -x^2+3 x+4 = 0

OpenStudy (anonymous):

x = -1 or, 4

OpenStudy (anonymous):

4x(x-8)-5x(x-7)=-4 4x^2 - 32x - 5x^2 + 35 = -4 -x^2 -32x + 35 = -4 0 = x^2 + 32x - 39 x^2 + 32x - 39 = 0 x = -16 - Sqrt[295] and x = -16 + Sqrt[295]